# High-dimensional sliced inverse regression with endogeneity

### v1

@@ Acknowledgments
$=E\left[E\left(\widehat{\mathbf{X}}\mid\mathbf{B}^{\top}\mathbf{X}\right)\mid y\right]$ (since $\varepsilon$ and $\widehat{\mathbf{X}}$ are independent)
=$\stackrel{{\scriptstyle(i)}}{{=}}\operatorname{\text{Cov}}(\widehat{\mathbf{X}},\mathbf{X})\mathbf{B}\left[\operatorname{\text{Var}}(\mathbf{B}^{\top}\mathbf{X})\right]^{-1}\mathbf{B}^{\top}E(\mathbf{X}\mid y),$
@@
S2.4 Proof of Lemma 2
Recall $\bm{\eta}$ and $\hat{\bm{\eta}}$ are the eigenvectors associated with $\bm{\Lambda}$ and $\widehat{\bm{\Lambda}}$, respectively. Hence, $\hat{\bm{\eta}}$ maximizes the function $^{\top}\hat{\bm{\Lambda}}\textbf{v}$ v subject to $\left\lVert\textbf{v}\right\rVert=1$, and
$\bm{\eta}^{\top}\hat{\bm{\Lambda}}\bm{\eta}\leq\hat{\bm{\eta}}^{\top}\hat{\bm{\Lambda}}\hat{\bm{\eta}}.$
[…]
For the left hand side of (13), since $\bm{\Lambda}=\lambda\bm{\eta}\bm{\eta}^{\top}$ and $\bm{\eta}^{\top}\bm{\eta}=1$, then we have $\langle\bm{\Lambda},\bm{\eta}\bm{\eta}^{\top}-\hat{\bm{\eta}}\hat{\bm{\eta}}^{\top}\rangle=\lambda(1-\hat{\rho}^{2})$.
Turning to the right hand side of (13), applying Hölder’s inequality we obtain
$\langle\hat{\bm{\Lambda}}-\bm{\Lambda},\bm{\eta}\bm{\eta}^{\top}-\hat{\bm{\eta}}\hat{\bm{\eta}}^{\top}\rangle\leq\left\lVert\hat{\bm{\Lambda}}-\bm{\Lambda}\right\rVert_{\max}\left\lVert\bm{\eta}\bm{\eta}^{\top}-\hat{\bm{\eta}}\hat{\bm{\eta}}^{\top}\right\rVert_{1}\leq\sqrt{2}\left\lVert\hat{\bm{\Lambda}}-\bm{\Lambda}\right\rVert_{\max},$
where for any matrix $\mathbf{A}$, we let $\left\lVert\mathbf{A}\right\rVert_{\max}$ denote its elementwise maximum norm, and the second inequality above follows from $\left\lVert\bm{\eta}\bm{\eta}^{\top}-\hat{\bm{\eta}}\hat{\bm{\eta}}^{\top}\right\rVert_{1}\leq\sqrt{\left\lVert\bm{\eta}\right\rVert_{2}^{2}+\left\lVert\hat{\bm{\eta}}\right\rVert_{2}^{2}}=\sqrt{2}$. Section S2.5 studies the properties of $\widehat{\bm{\Lambda}}$; in particular, it establishes that there exist positive constants $C$ and $C^{\prime}$ such that
$\left\lVert\hat{\bm{\Lambda}}-\bm{\Lambda}\right\rVert_{\max}\leq CL\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\}$
with probability at least $1-\exp(-C^{\prime}\log q)-\exp(-C^{\prime\prime}\log pq)$. With the same probability then, we obtain
$\hat{\rho}^{2}\geq 1-\frac{CL}{\lambda}\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\},$
as required.
S2.5 Properties of $\hat{\bm{\Lambda}}$
The developments below follow closely the work of Tan et al., (2018). Note
$\bm{\Lambda}=\operatorname{\text{Cov}}\left\{E\left(\widehat{\mathbf{X}}\mid y\right)\right\}=\bm{\Gamma}^{\top}\operatorname{\text{Cov}}\left\{E\left({\mathbf{Z}}\mid y\right)\right\}\bm{\Gamma}=\bm{\Gamma}^{\top}\left(\bm{\Sigma}_{\mathbf{Z}}-{\mathbf{\Xi}}\right)\bm{\Gamma},$
where $\mathbf{\Xi}=E\left\{\operatorname{\text{Cov}}\left({\mathbf{Z}}\mid y\right)\right\}$. Similarly, we can write $\widehat{\bm{\Lambda}}=\widehat{\bm{\Gamma}}^{\top}(\widehat{\bm{\Sigma}}_{\mathbf{Z}}-\widehat{\mathbf{\Xi}})\widehat{\bm{\Gamma}}$, with
$\widehat{\mathbf{\Xi}}=\dfrac{1}{H(c-1)}\mathcal{Z}^{\top}(\mathbf{I}_{H}\otimes\mathbf{P}_{c})\mathcal{Z},$ (14)
with $\mathbf{P}_{c}=\mathbf{I}_{c}-\mathbf{e}_{c}\mathbf{e}_{c}^{\top}$, $\mathbf{e}_{c}=(1,\ldots,1)^{\top}$ being the $c\times 1$ vector of ones, and $\mathbf{I}_{c}$ denoting the $c\times c$ identity matrix. Define $\mathbf{Q}_{c}=\mathbf{P}_{c}-\left(1-c^{-1}\right)\mathbf{I}_{c}=c^{-1}\left(\mathbf{I}_{c}-\mathbf{e}_{c}\mathbf{e}_{c}^{T}\right)$, and let $\bm{\Omega}=\bm{\Sigma}_{\mathbf{Z}}-\mathbf{\Xi}$, and $\widehat{\bm{\Omega}}=\widehat{\bm{\Sigma}}_{\mathbf{Z}}-\widehat{\mathbf{\Xi}}$.
We now bound the elementwise maximum norm $\left\lVert\widehat{\bm{\Lambda}}-\bm{\Lambda}\right\rVert_{\max}$. By the triangle inequality, for any $j,k=1,\ldots,q$, we have
@@
$\leq\|\hat{\bm{\gamma}}_{j}-\bm{\gamma}_{j}\|_{1}\left\lVert\bm{\Omega}\right\rVert_{\max}\left\lVert\bm{\gamma}_{k}\right\rVert_{1}+\|\hat{\bm{\gamma}}_{j}\|_{1}\|\bm{\Omega}-\widehat{\bm{\Omega}}\|_{\max}\|\bm{\gamma}_{k}\|_{1}+\left\lVert\widehat{\bm{\Omega}}\right\rVert_{\max}\left\lVert\hat{\bm{\gamma}}_{j}\right\rVert_{1}\|\hat{\bm{\gamma}}_{k}-\bm{\gamma}_{k}\|_{1}$
@@
From Theorem 1 in the main text we have $\|\hat{\bm{\gamma}}_{j}-\bm{\gamma}_{j}\|_{1}=C\sigma_{\max}r\sqrt{\log(pq)/n}$ with probability at least $1-\exp(-C^{\prime}\log pq)$ for any $j\in 1,\ldots,q$. Furthermore by Condition 4.2 in the main text, we have $\left\lVert\bm{\gamma}_{k}\right\rVert_{1}\leq L$ and $\|\bm{\Omega}\|_{\max}\leq C$. Therefore, obtain
@@
The proof for Lemma 4and other technical results in this section are deferred to section S3. Combining Lemma 1 and union bound then, we obtain
@@
$\left\lVert\widehat{\bm{\Omega}}\right\rVert_{\max}\leq\left\lVert\widehat{\bm{\Omega}}-\bm{\Omega}\right\rVert_{\max}+\left\lVert{\bm{\Omega}}\right\rVert_{\max}\leq C,$
@@
$\left\lVert\widehat{\bm{\Lambda}}-\bm{\Lambda}\right\rVert_{\max}=\max_{j,k}|\bm{\gamma}_{j}^{\top}\bm{\Omega}\bm{\gamma}_{k}-\hat{\bm{\gamma}}_{j}^{\top}\widehat{\bm{\Omega}}\hat{\bm{\gamma}}_{k}|\leq CL\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\},$ (19)
with a probability at least $1-\exp(-C^{\prime}\log pq)-\exp(-C^{\prime\prime}\log pq)$.
S2.6 Proof of Lemma 3
@@
$\dfrac{1}{2n}\|\tilde{\mathbf{y}}-\hat{\mathcal{X}}\hat{\bm{\beta}}\|_{2}^{2}+\mu_{2}\|\hat{\bm{\beta}}\|_{1}\leq\dfrac{1}{2n}\|\tilde{\mathbf{y}}-\hat{\mathcal{X}}{\tilde{\bm{\beta}}}\|_{2}^{2}+\mu_{2}\|{\tilde{\bm{\beta}}}\|_{1}.$ (20)
@@
$\dfrac{1}{2n}\|\tilde{\mathbf{y}}-\hat{\mathcal{X}}\tilde{\bm{\beta}}-\hat{\mathcal{X}}\bm{\delta}\|_{2}^{2}+\mu_{2}\|\bm{\delta}+\tilde{\bm{\beta}}\|_{1}\leq\dfrac{1}{2n}\|\tilde{\mathbf{y}}-\hat{\mathcal{X}}{\tilde{\bm{\beta}}}\|_{2}^{2}+\mu_{2}\|{\tilde{\bm{\beta}}}\|_{1}.$
@@
$\dfrac{1}{2n}\|\hat{\mathcal{X}}\bm{\delta}\|_{2}^{2}-\dfrac{1}{n}\bm{\delta}^{\top}\hat{\mathcal{X}}^{\top}(\tilde{\mathbf{y}}-\hat{\mathcal{X}}\tilde{\bm{\beta}})\leq\mu_{2}\left(\|\tilde{\bm{\beta}}\|_{1}-\|\tilde{\bm{\beta}}+\bm{\delta}\|_{1}\right).$ (21)
By definition, we have $\hat{\bm{\eta}}=n^{-1}\hat{\mathcal{X}}^{\top}\tilde{\mathbf{y}}$, and so (21) is equivalent to
@@
$\bm{\delta}^{\top}\left(\hat{\bm{\eta}}-\dfrac{1}{n}\hat{\mathcal{X}}^{\top}\hat{\mathcal{X}}\tilde{\bm{\beta}}\right)$ $\leq\|\bm{\delta}\|_{1}\left\|\hat{\bm{\eta}}-\dfrac{1}{n}\hat{\mathcal{X}}^{\top}\hat{\mathcal{X}}\tilde{\bm{\beta}}\right\|_{\infty}\leq\dfrac{1}{2}\mu_{2}\|\bm{\delta}\|_{1}.$
@@
$\left\lVert\tilde{{\bm{\beta}}}\right\rVert_{1}-\left\lVert\tilde{\bm{\beta}}+\bm{\delta}\right\rVert_{1}=\left\lVert\tilde{{\bm{\beta}}}_{S}\right\rVert_{1}-\left\lVert\tilde{{\bm{\beta}}}_{S}+{\bm{\delta}}_{S}\right\rVert_{1}-\left\lVert\bm{\delta}_{S^{c}}\right\rVert_{1}\leq\left\lVert{\bm{\delta}}_{S}\right\rVert_{1}-\left\lVert{\bm{\delta}}_{S^{c}}\right\rVert_{1}.$
Since $\left\lVert{\bm{\delta}}\right\rVert_{1}=\left\lVert{\bm{\delta}}_{S}\right\rVert_{1}+\left\lVert{\bm{\delta}}_{S^{c}}\right\rVert_{1}$, then equation (22) further leads to
00 0 $\leq\dfrac{1}{2}\left\lVert{\bm{\delta}}\right\rVert_{1}+\left\lVert\tilde{{\bm{\beta}}}\right\rVert_{1}-\left\lVert\tilde{\bm{\beta}}+\bm{\delta}\right\rVert_{1}$
$\stackrel{{\scriptstyle(iv)}}{{\leq}}\dfrac{1}{2}{\left\lVert{\bm{\delta}}_{S}\right\rVert_{1}+\dfrac{1}{2}\left\lVert{\bm{\delta}}_{S^{c}}\right\rVert_{1}}+\left\lVert{\bm{\delta}}_{S}\right\rVert_{1}-\left\lVert{\bm{\delta}}_{S^{c}}\right\rVert_{1}=\dfrac{3}{2}\left\lVert{\bm{\delta}}_{S}\right\rVert_{1}-\dfrac{1}{2}\left\lVert{\bm{\delta}}_{S^{c}}\right\rVert_{1}$
$\stackrel{{\scriptstyle(v)}}{{\leq}}\dfrac{3}{2}\left\lVert{\bm{\delta}}_{S}\right\rVert_{1}.$
It follows from step (iv) that $\left\lVert{\bm{\delta}}_{S^{c}}\right\rVert_{1}\leq 3\left\lVert{\bm{\delta}}_{S}\right\rVert_{1}$, or $\bm{\delta}\in\mathcal{C}(S,3)$.
@@
Assume Conditions 4.1-4.2, 4.5 and the requirements of Theorem 1 are satisfied. Then there exists a constant $\kappa_{2}>0$ such that $\left\lVert\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}^{1/2}\bm{\delta}\right\rVert_{2}^{2}\geq\kappa_{2}\left\lVert\bm{\delta}\right\rVert_{2}^{2}$, for any $\delta\in\mathcal{C}(S,3)$, with probability at least $1-\exp(-C^{\prime}\log q)-\exp(-C^{\prime\prime}\log pq)$, for constants $C^{\prime}$ and $C^{\prime\prime}$.
@@
$\begin{split}\dfrac{1}{2}\kappa_{2}\left\lVert\bm{\delta}\right\rVert_{2}^{2}{\leq}\dfrac{1}{2n}\left\lVert\hat{\mathcal{X}}{\bm{\delta}}\right\rVert_{2}^{2}&{\leq}\dfrac{1}{2}\mu_{2}\left\lVert{\bm{\delta}}\right\rVert_{1}+\mu_{2}\left(\left\lVert\tilde{{\bm{\beta}}}\right\rVert_{1}-\left\lVert\hat{{\bm{\beta}}}\right\rVert_{1}\right)\\ &{\leq}\dfrac{3}{2}\mu_{2}\left\lVert\bm{\delta}_{S}\right\rVert_{1}{\leq}\dfrac{3}{2}\mu_{2}\sqrt{s}\left\lVert\bm{\delta}_{S}\right\rVert_{2}\leq\dfrac{3}{2}\mu_{2}\sqrt{s}\left\lVert\bm{\delta}\right\rVert_{2}.\end{split}$
@@
$\left\lVert\hat{\bm{\beta}}-\tilde{\bm{\beta}}\right\rVert_{2}=\left\lVert\bm{\delta}\right\rVert_{2}\leq\dfrac{3}{\kappa_{2}}\mu_{2}\sqrt{s},$
@@
S2.7 Proof of Theorem 2
Recall that $\tilde{\bm{\beta}}=\xi^{-1}\bm{\beta}(\hat{\bm{\eta}}^{\top}\bm{\eta})$ with $\xi=\left\lVert\Sigma_{\widehat{\mathbf{X}}}\bm{\beta}\right\rVert_{2}>0$. By Lemma 2, the term $\hat{\bm{\eta}}^{\top}\bm{\eta}$ is non-zero with high probability. Hence, $P(\bm{\beta})=P(\hat{\bm{\beta}})$ with the same probability. Next, by the Cauchy-Schwartz inequality we have

### v2

@@ B.3 Proof of Lemma 1
$=E\left[E\left(\widehat{\mathbf{X}}\mid\mathbf{B}^{\top}\mathbf{X}\right)\mid y\right]~\quad~~~\text{(since $\varepsilon$ and $\widehat{\mathbf{X}}$ are independent)}$
$\stackrel{{\scriptstyle(i)}}{{=}}\operatorname{\text{cov}}(\widehat{\mathbf{X}},\mathbf{X})\mathbf{B}\left[\operatorname{\text{Var}}(\mathbf{B}^{\top}\mathbf{X})\right]^{-1}\mathbf{B}^{\top}E(\mathbf{X}\mid y),$
@@
B.4 Properties of $\hat{\bm{\Lambda}}$
The developments below closely follow the work of Tan et al., (2018). Note
$\bm{\Lambda}=\operatorname{\text{cov}}\left\{E\left(\widehat{\mathbf{X}}\mid y\right)\right\}=\bm{\Gamma}^{\top}\operatorname{\text{cov}}\left\{E\left({\mathbf{Z}}\mid y\right)\right\}\bm{\Gamma}=\bm{\Gamma}^{\top}\left(\bm{\Sigma}_{\mathbf{Z}}-{\bm{\Xi}}\right)\bm{\Gamma},$
where $\bm{\Xi}=E\left\{\operatorname{\text{cov}}\left({\mathbf{Z}}\mid y\right)\right\}$. Similarly, we can write $\widehat{\bm{\Lambda}}=\widehat{\bm{\Gamma}}^{\top}(\widehat{\bm{\Sigma}}_{\mathbf{Z}}-\widehat{\bm{\Xi}})\widehat{\bm{\Gamma}}$, with
$\widehat{\bm{\Xi}}=\dfrac{1}{H(c-1)}\mathcal{Z}^{\top}(\mathbf{I}_{H}\otimes\mathbf{P}_{c})\mathcal{Z},$ (14)
with $\mathbf{P}_{c}=\mathbf{I}_{c}-\bm{e}_{c}\bm{e}_{c}^{\top}$, $\mathbf{e}_{c}=(1,\ldots,1)^{\top}$ being the $c\times 1$ vector of ones, and $\mathbf{I}_{c}$ denoting the $c\times c$ identity matrix. Define $\mathbf{Q}_{c}=\mathbf{P}_{c}-\left(1-c^{-1}\right)\mathbf{I}_{c}=c^{-1}\left(\mathbf{I}_{c}-\mathbf{e}_{c}\mathbf{e}_{c}^{T}\right)$, and let $\bm{\Omega}=\bm{\Sigma}_{\mathbf{Z}}-\bm{\Xi}$, and $\widehat{\bm{\Omega}}=\widehat{\bm{\Sigma}}_{\mathbf{Z}}-\widehat{\bm{\Xi}}$.
We now bound the elementwise maximum norm $\lVert\widehat{\bm{\Lambda}}-\bm{\Lambda}\rVert_{\max}$. By the triangle inequality, for any $j,k=1,\ldots,q$, we have
@@ B.4 Properties of $\hat{\bm{\Lambda}}$
$\leq\|\hat{\bm{\gamma}}_{j}-\bm{\gamma}_{j}\|_{1}\lVert\bm{\Omega}\rVert_{\max}\lVert\bm{\gamma}_{k}\rVert_{1}+\|\hat{\bm{\gamma}}_{j}\|_{1}\|\bm{\Omega}-\widehat{\bm{\Omega}}\|_{\max}\|\bm{\gamma}_{k}\|_{1}+\lVert\widehat{\bm{\Omega}}\rVert_{\max}\lVert\hat{\bm{\gamma}}_{j}\rVert_{1}\|\hat{\bm{\gamma}}_{k}-\bm{\gamma}_{k}\|_{1}$
@@
From Theorem 1 in the main text we have $\|\hat{\bm{\gamma}}_{j}-\bm{\gamma}_{j}\|_{1}=C\sigma_{\max}r\sqrt{\log(pq)/n}$ with probability at least $1-\exp(-C^{\prime}\log pq)$ for any $j\in 1,\ldots,q$. Furthermore by Condition 4.2 in the main text, we have $\lVert\bm{\gamma}_{k}\rVert_{1}\leq L$ and $\|\bm{\Omega}\|_{\max}\leq C$. Therefore, obtain
@@
The proof for Lemma 3and other technical results in this section are deferred to section C. Combining Lemma 1 and union bound then, we obtain
@@
$\lVert\widehat{\bm{\Omega}}\rVert_{\max}\leq\lVert\widehat{\bm{\Omega}}-\bm{\Omega}\rVert_{\max}+\lVert{\bm{\Omega}}\rVert_{\max}\leq C,$
@@
$\lVert\widehat{\bm{\Lambda}}-\bm{\Lambda}\rVert_{\max}=\max_{j,k}|\bm{\gamma}_{j}^{\top}\bm{\Omega}\bm{\gamma}_{k}-\hat{\bm{\gamma}}_{j}^{\top}\widehat{\bm{\Omega}}\hat{\bm{\gamma}}_{k}|\leq CL\left\{L\sqrt{\dfrac{\log q}{n}}+\sigma_{\max}r\sqrt{\dfrac{\log pq}{n}}\right\},$ (19)
Eigenvalue and eigenvector analyses for $\hat{\bm{\Lambda}}$
Recall that at the population level ${\bm{\Lambda}}={\bm{\Gamma}}^{\top}{\bm{\Omega}}\bm{\Gamma}$, where $\bm{\Omega}$ has rank $d$. Let $\bm{\eta}_{j}(\bm{\Omega})$ be the eigenvector associated with the $j$th largest eigenvalue $\lambda_{j}(\bm{\Omega})$ of $\bm{\Omega}$ for $j=1,\ldots,d$, and we also write $\lambda(\bm{\Omega})=\lambda_{1}(\bm{\Omega})$ and $\bm{\eta}(\bm{\Omega})=\bm{\eta}_{1}(\bm{\Omega})$ (i.e we suppress the subscript 1 when we refer to the largest eigenvalue). When $d=1$, the (only) non-zero eigenvalue of $\bm{\Lambda}$ is $\lambda(\bm{\Lambda})=\lambda(\bm{\Omega})\|\bm{\Gamma}^{\top}\bm{\eta}(\bm{\Omega})\|_{2}$, but if $d>1$, there are generally no closed form for the eigenvalues of $\bm{\Lambda}$. Nevertheless, because $\bm{\Omega}$ has rank $d\leq\min(p,q)$, applying the Weyl’s inequality yields
$\nu_{\min}^{2}(\bm{\Gamma})\lambda_{j}(\bm{\Omega})\leq\lambda_{j}(\bm{\Lambda})\leq\nu_{\max}^{2}(\bm{\Gamma})\lambda_{j}(\bm{\Omega}),$ (20)
for $j=1,\ldots,d$, where $\nu_{\min}(\bm{\Gamma})$ and $\nu_{\max}(\bm{\Gamma})$ denote the minimum singular value and maximum singular value for $\bm{\Gamma}$. Similar inequalities hold for $\widehat{\bm{\Lambda}}$, respectively. Furthermore, since conditions 4.4–4.6 are imposed, the results from Lin et al., (2019) are applied to various quantities related to $\widehat{\bm{\Omega}}$, which we state in Proposition 3 below.
Proposition 3.
If $n\lambda(\bm{\Omega})=q^{\alpha}$ for some $\alpha>1/2$, there exist positive constants $C_{1}$ and $C_{2}$ such that, with a probability of at least $1-\exp(C^{\prime}\log(q))$
i) for $j=1,\ldots.,d$, one has
$\|P_{\bm{\Omega}}{\bm{\eta}}_{j}(\widehat{\bm{\Omega}})\|_{2}\geq C_{1}\sqrt{\frac{\lambda(\bm{\Omega})}{{\lambda}_{j}(\widehat{\bm{\Omega}})}}~~$,
ii) for $j=d+1,\ldots,H$, one has
$\|P_{\bm{\Omega}}{\bm{\eta}}_{j}(\widehat{\bm{\Omega}})\|_{2}\leq C_{2}\frac{\sqrt{q\log(q)}}{n\lambda(\bm{\Omega})}\sqrt{\frac{\lambda(\bm{\Omega})}{{\lambda_{j}}(\widehat{\bm{\Omega}})}}$.
With these results, combining with the results from Theorem 1, we then have for $j=1,\ldots,d$,
$\dfrac{\lambda(\bm{\Lambda})}{\lambda_{j}(\bm{\widehat{\Lambda}})}\leq\dfrac{\lambda(\bm{\Omega})}{\lambda_{j}(\widehat{\bm{\Omega}})}\times\dfrac{\nu_{\max}(\bm{\Gamma})}{\nu_{\min}(\widehat{\bm{\Gamma}})}\leq C,$ (21)
with probability of at least $1-\exp(-C^{\prime}\log q)-\exp(-C^{\prime\prime}\log pq)$.
B.5 Proof of Lemma 3
@@
$\dfrac{1}{2n}\|\tilde{{y}}-\hat{\mathcal{X}}\hat{\bm{\beta}}\|_{2}^{2}+\mu_{2}\|\hat{\bm{\beta}}\|_{1}\leq\dfrac{1}{2n}\|\tilde{{y}}-\hat{\mathcal{X}}{\tilde{\bm{\beta}}}\|_{2}^{2}+\mu_{2}\|{\tilde{\bm{\beta}}}\|_{1}.$ (22)
@@
$\dfrac{1}{2n}\|\tilde{{y}}-\hat{\mathcal{X}}\tilde{\bm{\beta}}-\hat{\mathcal{X}}\bm{\delta}\|_{2}^{2}+\mu_{2}\|\bm{\delta}+\tilde{\bm{\beta}}\|_{1}\leq\dfrac{1}{2n}\|\tilde{{y}}-\hat{\mathcal{X}}{\tilde{\bm{\beta}}}\|_{2}^{2}+\mu_{2}\|{\tilde{\bm{\beta}}}\|_{1}.$
@@
$\dfrac{1}{2n}\|\hat{\mathcal{X}}\bm{\delta}\|_{2}^{2}-\dfrac{1}{n}\bm{\delta}^{\top}\hat{\mathcal{X}}^{\top}(\tilde{{y}}-\hat{\mathcal{X}}\tilde{\bm{\beta}})\leq\mu_{2}\left(\|\tilde{\bm{\beta}}\|_{1}-\|\tilde{\bm{\beta}}+\bm{\delta}\|_{1}\right).$ (23)
By definition, we have $\hat{\bm{\eta}}=n^{-1}\hat{\mathcal{X}}^{\top}\tilde{y}$, and so (23) is equivalent to
@@
[…]
@@
$\lVert\tilde{{\bm{\beta}}}\rVert_{1}-\lVert\tilde{\bm{\beta}}+\bm{\delta}\rVert_{1}=\lVert\tilde{{\bm{\beta}}}_{S}\rVert_{1}-\lVert\tilde{{\bm{\beta}}}_{S}+{\bm{\delta}}_{S}\rVert_{1}-\lVert\bm{\delta}_{S^{c}}\rVert_{1}\leq\lVert{\bm{\delta}}_{S}\rVert_{1}-\lVert{\bm{\delta}}_{S^{c}}\rVert_{1}.$
Since $\lVert{\bm{\delta}}\rVert_{1}=\lVert{\bm{\delta}}_{S}\rVert_{1}+\lVert{\bm{\delta}}_{S^{c}}\rVert_{1}$, then equation (24) further leads to
$0$ $\leq\dfrac{1}{2}\lVert{\bm{\delta}}\rVert_{1}+\lVert\tilde{{\bm{\beta}}}\rVert_{1}-\lVert\tilde{\bm{\beta}}+\bm{\delta}\rVert_{1}$
$\stackrel{{\scriptstyle(iv)}}{{\leq}}\dfrac{1}{2}{\lVert{\bm{\delta}}_{S}\rVert_{1}+\dfrac{1}{2}\lVert{\bm{\delta}}_{S^{c}}\rVert_{1}}+\lVert{\bm{\delta}}_{S}\rVert_{1}-\lVert{\bm{\delta}}_{S^{c}}\rVert_{1}=\dfrac{3}{2}\lVert{\bm{\delta}}_{S}\rVert_{1}-\dfrac{1}{2}\lVert{\bm{\delta}}_{S^{c}}\rVert_{1}$
$\stackrel{{\scriptstyle(v)}}{{\leq}}\dfrac{3}{2}\lVert{\bm{\delta}}_{S}\rVert_{1}.$
It follows from step (iv) that $\lVert{\bm{\delta}}_{S^{c}}\rVert_{1}\leq 3\lVert{\bm{\delta}}_{S}\rVert_{1}$, or $\bm{\delta}\in\mathcal{C}(S,3)$.
@@
Assume Conditions 4.1-4.2, 4.5 and the requirements of Theorem 1 are satisfied. Then there exists a constant $\kappa_{2}>0$ such that $\lVert\widehat{\bm{\Sigma}}_{\widehat{\mathbf{X}}}^{1/2}\bm{\delta}\rVert_{2}^{2}\geq\kappa_{2}\lVert\bm{\delta}\rVert_{2}^{2}$, for any $\delta\in\mathcal{C}(S,3)$, with probability at least $1-\exp(-C^{\prime}\log q)-\exp(-C^{\prime\prime}\log pq)$, for constants $C^{\prime}$ and $C^{\prime\prime}$.
@@
$\begin{split}\dfrac{1}{2}\kappa_{2}\lVert\bm{\delta}\rVert_{2}^{2}{\leq}\dfrac{1}{2n}\lVert\hat{\mathcal{X}}{\bm{\delta}}\rVert_{2}^{2}&{\leq}\dfrac{1}{2}\mu_{2}\lVert{\bm{\delta}}\rVert_{1}+\mu_{2}\left(\lVert\tilde{{\bm{\beta}}}\rVert_{1}-\lVert\hat{{\bm{\beta}}}\rVert_{1}\right)\\ &{\leq}\dfrac{3}{2}\mu_{2}\lVert\bm{\delta}_{S}\rVert_{1}{\leq}\dfrac{3}{2}\mu_{2}\sqrt{s}\lVert\bm{\delta}_{S}\rVert_{2}\leq\dfrac{3}{2}\mu_{2}\sqrt{s}\lVert\bm{\delta}\rVert_{2}.\end{split}$
@@
$\lVert\hat{\bm{\beta}}-\tilde{\bm{\beta}}\rVert_{2}=\lVert\bm{\delta}\rVert_{2}\leq\dfrac{3}{\kappa_{2}}\mu_{2}\sqrt{s},$
@@
B.6 Proof of Theorem 2
[…]